# Learning Value of Information towards Joint Communication and Control in 6G V2X

### v1

@@ V-A Utility-based VoI
  • EVoMI $\Xi_{i}^{\rm M}=J_{\pi_{i}^{*}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{*}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the optimal policies $\pi_{i}^{*}$ and $\tilde{\pi}_{i}^{*}$ for the SSDP in Case 2 and augmented-state SSDP in Case 4, respectively.
  • IVoMI $\xi_{i,k}^{\rm M}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\pi_{i}^{*}(\cdot|S_{i,k})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the augmented-state SSDP in Case 4, and $u_{i,k}\sim\pi_{i}^{*}(\cdot|S_{i,k})$ is the control input of vehicle $i$ that is sampled according to the optimal policy $\pi_{i}^{*}(\cdot|S_{i,k})$ of the SSDP in Case 2.
  • EVoII $\Xi_{i}^{\rm I}=J_{\pi_{i}^{\rm d*}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{\rm d*}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the optimal policies $\pi_{i}^{\rm d*}$ and $\tilde{\pi}_{i}^{*}$ for the Random Delay SSDP in Case 5 and augmented-state SSDP in Case 4, respectively.
  • IVoII $\xi_{i,k}^{\rm I}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\tilde{\pi}_{i}^{\rm d*}(\cdot|\tilde{S}_{i,k}^{\rm d})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the augmented-state SSDP in Case 4, and $u_{i,k}\sim\tilde{\pi}_{i}^{\rm d*}(\cdot|\tilde{S}_{i,k}^{\rm d})$ is the control input of vehicle $i$ that is sampled according to the optimal policy $\tilde{\pi}_{i}^{\rm d*}(\cdot|\tilde{S}_{i,k}^{\rm d})$ of the Random Delay SSDP in Case 5.
@@ V-B Information-Theory-based VoI
According to Lemma 1, given the SSDP in Definition 1, a piece of information $I_{k}$ can improve decision making only if it is part of the exogenous information $W_{k}$ or helps in predicting $W_{k}$. Furthermore, as per Definition 1, $W_{k}$ must either affect the transition of state $S_{k}$ and/or the reward $R_{k+1}$. Thus, it follows intuitively that the value of $I_{k}$ is related to its impact on both the state transition and the reward. The VoI can therefore be defined by quantifying, ”How much better could we predict the next state $S_{k+1}$ or the reward $R_{k+1}$ if we included $I_{k}$ in the augmented-state $\tilde{S}_{k}$, compared to if we did not?”
@@
Compared to the utility-based VoI, the ITVoI is generally easier to calculate when the model of the environment, i.e., the state transition probabilities and the reward function, is known. This advantage arises because ITVoI does not require deriving the optimal policies of the SSDPs with state $S_{k}$ and augmented state $\tilde{S}_{k}$. However, there are several significant limitations to ITVoI. Firstly, it requires the knowledge of the model of the environment, which might not be accurately available in many practical tasks. Secondly, among the four types of utility-based VoI, the ITVoI is essentially equivalent to the EVoMI, as both aim to predict the benefit of having additional information $I_{k}$ before making decisions in terms of long-term average performance. However, ITVoI does not cover the other three types of utility-based VoI, i.e., IVoMI, EVoII, and IVoII, thereby limiting its applicability. Finally, the utility-based VoI captures the ultimate objective of information usage, which is to improve the control performance. Meanwhile, the ITVoI measures the merit of information for improving some intermediate outcomes that could contribute to the enhancement of the ultimate objective. Therefore, utility-based VoI is more intuitive compared to ITVoI. In the next section, we will focus on the estimation of utility-based VoI.
@@ VI VoI Estimation using DRL
To obtain either EVoI or IVoI, we have to derive both policies $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}_{k}$ in Definition 7 and Definition 8. This can be achieved by using the DRL algorithms to train the control agent. The inferior policy $\pi^{\rm{inf}}$ is learned in the environment with partial/imperfect observations $S^{\rm{inf}}_{k}$, while the superior policy $\pi^{\rm{sup}}$ is learned in the environment with full/perfect observations $S^{\rm{sup}}_{k}$.
@@ VI-B IVoI Estimation
The IVoI corresponds to the advantage function of the superior policy $\pi^{\rm{sup}}$ with full/perfect observations $S^{\rm{sup}}_{k}$, where various methods can be used to estimate the advantage function of a policy in RL.
@@ VI-B1 Method A
By the definition of advantage function in (14), we can train two neural networks (NNs) to represent the value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$ and Q function $q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k})$, respectively. The IVoI can be derived simply as their difference
@@
However, it is a little cumbersome to train two NNs. Moreover, if the NN that represents the Q function $q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k})$ is trained during learning of the superior policy $\pi^{\rm{sup}}$, it might provide a less accurate estimation of the Q values for those actions that are favored by the inferior policy $\pi^{\rm{inf}}$ but seldomly chosen by $\pi^{\rm{sup}}$.
VI-B2 Method B
To solve the above issues, another method is to only train one NN to represent the value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$, and estimate the advantage function using the Temporal Difference (TD) error
@@ VI-B2 Method B
In this case, the IVoI can be estimated for each control interval based on the reward $R_{k+1}$ as a result of implementing action $a_{k}^{\rm{inf}}\sim(a_{k}|S^{\rm{inf}}_{k})$, as well as the value function $v_{\pi^{\rm{sup}}}$ of the optimal control policy for the current state $S^{\rm{sup}}_{k}$ and the next state $S^{\rm{sup}}_{k+1}$. Since we have
$=\mathrm{E}_{S^{\rm{sup}}_{k+1}}[\delta_{\pi^{\rm{inf}},k}],$ (31)
$\delta_{\pi^{\rm{inf}},k}$ is an unbiased estimate of the IVoI if the NN can accurately represent the true value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$. This method of estimation is subject to noise. However, the variance remains relatively low, as the expectation in (VI-B2) is taken only with respect to the next state $S^{\rm{sup}}_{k+1}$.
@@ VII-B3 Across Multiple Control Intervals - When and What to Communicate
  • Every Communication Interval - How to Communicate: The communication agent makes RRM decisions at each communication interval $(k,t)$. In order to enhance spectrum utilization, we consider that $L$ V2V links share the uplink resources orthogonally allocated to $M$ V2I links. Without loss of generality, we assume that every V2I link $m\in\mathcal{M}=\{0,\cdots,M-1\}$ is pre-assigned sub-channel $m$ with constant transmit power $P^{\rm I}_{m}$. Moreover, one or more V2V links $i-1\in\mathcal{L}=\{1,3,5\cdots,2L-1\}$ can reuse the sub-channels of the V2I links for CAM transmission. We use the binary allocation indicator $\theta_{i-1,m,(k,t)}\in\{0,1\}$ to indicate whether V2V link $i-1$ occupies sub-channel $m$ at communication interval $(k,t)$ or not. Moreover, we consider that each V2V link $i-1$ occupies at most one sub-channel, i.e., $\sum_{m=0}^{M-1}\theta_{i-1,m,(k,t)}\leq 1$. Let $P^{\rm V}_{i-1,m,(k,t)}$ denote the transmit power of V2V link $i-1$ over the sub-channel $m$ at communication interval $(k,t)$. Therefore, each communication agent $i-1$ has to make the subchannel allocation $\{\theta_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ and power allocation $\{P^{\rm V}_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ decisions at each communication interval $(k,t)$.
  • Every Control Interval - When and What to Communicate: Let $\phi_{i-1,k}\in\{0,1\}$ be the binary transmission indicator to indicate whether the sampled CAM $acc_{i-1,k}$ will be transmitted to the follower $i$ or not. At each control interval $k$, each communication agent $i-1$ can make decisions on $\phi_{i-1,k}$.
  • Across Multiple Control Intervals - When and What to Communicate: Each communication agent $i-1$ makes a one-time decision on whether it will transmit its acceleration or not to its follower $i$. Once the decision is made, the predecessor will always transmit $acc_{i-1,k}$ or not at each control interval $k$.

### v2

@@ V-A Utility-based VoI
  • EVoMI $\Xi_{i}^{\rm M}=J_{\pi_{i}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the acting policy $\pi_{i}$ based on partial observations $S_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k}]^{\mathrm{T}}$ (for the SSDP in Case 2) and the optimal policy $\tilde{\pi}_{i}^{*}$ based on full observations $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}$ (for the augmented-state SSDP in Case 4), respectively. The EVoMI quantifies the performance loss incurred when the following vehicle $i$ consistently fails to receive the acceleration $acc_{i-1,k}$ of the preceding vehicle $i-1$ at every control interval $k$ throughout the entire episode. Note that the performance refers to the expected cumulative reward defined in (13), while the reward function of the vehicle-following control problem, given in (2), depends on the position error, velocity error, and acceleration of vehicle $i$.
  • IVoMI $\xi_{i,k}^{\rm M}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\pi_{i}(\cdot|S_{i,k})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the optimal policy $\tilde{\pi}_{i}^{*}$, and $u_{i,k}\sim\pi_{i}(\cdot|S_{i,k})$ is the control input of vehicle $i$ that is sampled according to the acting policy $\pi_{i}(\cdot|S_{i,k})$ based on partial observation $S_{i,k}$. The IVoMI quantifies the performance loss incurred when the following vehicle $i$ fails to receive the acceleration $acc_{i-1,k}$ of the preceding vehicle $i-1$ at a specific control interval $k$, while consistently receiving this information during the rest of the episode. Consequently, the IVoMI is a function of the state action pair $\{\tilde{S}_{i,k},u_{i,k}\}$, where $\tilde{S}_{i,k}$ represents the full observation at control interval $k$, including $acc_{i-1,k}$. It measures the performance loss incurred when the control agent has access only to the partial observation (excluding $acc_{i-1,k}$) and selects action $u_{i,k}$ using the acting control policy. Unlike EVoI, IVoI distinguishes the value of knowing the predecessor’s acceleration based on both its specific values and the available system state information, such as position error, velocity error, and acceleration of the follower.
  • EVoII $\Xi_{i}^{\rm I}=J_{\pi_{i}^{\rm d}}-J_{\tilde{\pi}_{i}^{*}}$, where $J_{\pi_{i}^{\rm d}}$ and $J_{\tilde{\pi}_{i}^{*}}$ are the performance of the acting policy $\pi_{i}^{\rm d}$ based on imperfect (delayed) observations $\tilde{S}_{i,k}^{\rm d}=[x_{i,k-\tau_{i,k}},acc_{(i-1),k-\tau_{i,k}},\{u_{i,k^{\prime}}\}_{k^{\prime}=k-\tau_{\mathrm{max}}}^{k-1},\tau_{i,k}]^{\mathrm{T}}$ (for the Random Delay SSDP in Case 5) and the optimal policy $\tilde{\pi}_{i}^{*}$ based on perfect observations $\tilde{S}_{i,k}=[e_{pi,k},e_{vi,k},acc_{i,k},acc_{(i-1),k}]^{\mathrm{T}}$ (for the augmented-state SSDP in Case 4), respectively. The EVoII is similar to the EVoMI, with the only difference being that the performance loss results from the following vehicle $i$ receiving a delayed acceleration $acc_{i-1,k-\tau_{k}}$ from the predecessor $i-1$, rather than not receiving it at all.
  • IVoII $\xi_{i,k}^{\rm I}=A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})|_{u_{i,k}\sim\tilde{\pi}_{i}^{\rm d}(\cdot|\tilde{S}_{i,k}^{\rm d})}$, where $A_{\tilde{\pi}_{i}^{*}}(\tilde{S}_{i,k},u_{i,k})$ is the advantage function of the optimal policy $\tilde{\pi}_{i}^{*}$, and $u_{i,k}\sim\tilde{\pi}_{i}^{\rm d}(\cdot|\tilde{S}_{i,k}^{\rm d})$ is the control input of vehicle $i$ that is sampled according to the acting policy $\tilde{\pi}_{i}^{\rm d}(\cdot|\tilde{S}_{i,k}^{\rm d})$ based on delayed observation $\tilde{S}_{i,k}^{\rm d}$. The IVoII is similar to the IVoMI, in the same way that the EVoII parallels the EVoMI.
@@ V-B Information-Theory-based VoI
According to Lemma 1, given the SSDP in Definition 1, a piece of information $I_{k}$ can improve decision making only if it is part of the exogenous information $W_{k}$ or helps in predicting $W_{k}$. Furthermore, as per Definition 1, $W_{k}$ must either affect the transition of state $S_{k}$ and/or the reward $R_{k+1}$. Thus, it follows intuitively that the value of $I_{k}$ is related to its impact on both the state transition and the reward. VoI can thus be defined by quantifying, “How much better could we predict the next state $S_{k+1}$ or the reward $R_{k+1}$ if we included $I_{k}$ in the augmented-state $\tilde{S}_{k}$, compared to if we did not?” [35].
@@
We would like to emphasize that although KL divergence is used to quantify ITVoI, the ultimate objective remains to evaluate the impact of information on control performance, as established in Lemma 1. This contrasts with related concepts in information theory, such as information entropy, which aim to quantify uncertainty itself without regard to its effect on outcomes.
@@
Remark 5 (Comparison between utility-based VoI and ITVoI).
Compared to the utility-based VoI, the ITVoI is generally easier to estimate when the model of the environment, i.e., the state transition probabilities and the reward function, is known. This advantage arises because ITVoI does not require deriving the optimal policies of the SSDPs with state $S_{k}$ and augmented state $\tilde{S}_{k}$. However, there are several significant limitations to ITVoI. Firstly, it requires the knowledge of the environment model, which might not be accurately available in many practical tasks. Secondly, among the four types of utility-based VoI, the ITVoI is essentially equivalent to the EVoMI, as both aim to predict the benefit of having additional information $I_{k}$ before making decisions in terms of long-term average performance. However, ITVoI does not cover the other three types of utility-based VoI, i.e., IVoMI, EVoII, and IVoII, thereby limiting its applicability. Finally, the utility-based VoI captures the ultimate objective of information usage, which is to improve the control performance. Meanwhile, the ITVoI measures the merit of information for improving some intermediate outcomes that could contribute to the enhancement of the ultimate objective. Therefore, utility-based VoI is more intuitive compared to ITVoI.
@@ VI VoI Estimation using DRL
For the joint design of control and communication, VoI should be estimated by the control agents and subsequently utilized by the communication agents to guide the optimization of communication decisions. This section focuses on VoI estimation, while VoI-oriented optimization for communication agents is presented in Section VI.
In Section V, we have defined two types of VoI: utility-based VoI and ITVoI. When the environment model is known, the ITVoI can be directly derived from (26), (9), and (9). Otherwise, the distributions of next states and rewards must be estimated from data [35]. In the following, we focus on the estimation of utility-based VoI, including EVoI and IVoI.
To obtain either EVoI or IVoI, we require both policies $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}_{k}$ in Definition 7 and Definition 8. These policies can be derived using RL, dynamic programming, or optimal control algorithms. The superior policy $\pi^{\rm{sup}}$ is typically learned in a simulated environment with full/perfect observations, while the inferior policy $\pi^{\rm{inf}}$ can be learned in an environment with either partial/imperfect or full/perfect observations. In the latter case, the partial/imperfect observations are treated as full/perfect during policy execution.
@@ VI-B IVoI Estimation
The IVoI corresponds to the advantage function of the policy $\pi^{\rm{sup}}$ with full/perfect observations $S^{\rm{sup}}_{k}$, where various methods can be used to estimate the advantage function of a policy in RL.
@@ VI-B1 Method A
Based on the definition of IVoI in Definition 8, we can estimate it using the Monte Carlo method. Specifically, we begin by sampling initial states from their distribution and then generate multiple trajectories by executing three policies: $\pi^{\rm{inf}}$, $\pi^{\rm{sup}}$, and a mixed policy that randomly selects between $\pi^{\rm{inf}}$ and $\pi^{\rm{sup}}$ at each control interval. Each trajectory comprises a sequence of states, where each state contains a pair of full/perfect and partial/imperfect observations, $S^{\rm{sup}}_{k}$ and $S^{\rm{inf}}_{k}$, respectively, at control interval $k$.
From these trajectories, a subset of states is selected to form a rollout set. For each state in this set, we sample two actions: $a^{\rm{inf}}_{k}$ from the acting policy $\pi^{\rm{inf}}$ based on $S^{\rm{inf}}_{k}$, and $a^{\rm{sup}}_{k}$ from the optimal policy $\pi^{\rm{sup}}$ based on $S^{\rm{sup}}_{k}$, respectively. Two roll outs are then performed starting from the state action pairs $(S^{\rm{sup}}_{k},a^{\rm{inf}}_{k})$ and $(S^{\rm{sup}}_{k},a^{\rm{sup}}_{k})$, respectively, with the optimal policy $\pi^{\rm{sup}}$ consistently used thereafter.
The difference in cumulative rewards between the two roll outs is calculated for each state in the rollout set. These differences can serve as target values for training an advantage function estimator, e.g., a deep neural network (DNN), using supervised learning.
The main drawback of this method is that it requires a large amount of experience data to train the advantage estimator accurately.
VI-B2 Method B
Since the IVoI is essentially the advantage function of the optimal policy $\pi^{\rm{sup}}$ under full/perfect observations, it can be derived from the value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$ and Q function $q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k})$ of $\pi^{\rm{sup}}$ according to (14). Specifically, given $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$ and $q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},a_{k})$, the IVoI can be computed as their difference
@@ VI-B2 Method B
Using (VI-B2), we convert the advantage estimation problem into the task of estimating the value and Q functions of $\pi^{\rm{sup}}$. This can be achieved by training two separate estimators using a method similar to Method A. However, this approach is less appealing, as training two estimators is generally more cumbersome than training a single estimator for the advantage function directly.
Method B becomes more attractive in the case of deterministic policies, where the value function can be computed from the Q function as $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})=q_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k},\pi^{\rm{sup}}(S^{\rm{sup}}_{k}))$, so only the Q function needs to be estimated. Classical DRL algorithms such as deep Q-network (DQN) [36] and deep deterministic policy gradient (DDPG) [37] already learn Q functions during training. This implies that a Q function estimator is readily available when these algorithms are used to learn $\pi^{\rm{sup}}$. However, this estimator may be less accurate for actions preferred by the acting policy $\pi^{\rm{inf}}$ but rarely chosen by $\pi^{\rm{sup}}$. Therefore, it is advisable to fine-tune the Q function estimator using a Monte Carlo method similar to that used in Method A.
VI-B3 Method C
Another method is to only train a single estimator for the value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$ and estimate the advantage function using the Temporal Difference (TD) error
@@ VI-B3 Method C
In this case, the IVoI can be estimated at each control interval based on the reward $R_{k+1}$, resulting from executing the action $a_{k}^{\rm{inf}}\sim(\cdotp|S^{\rm{inf}}_{k})$, and the value function $v_{\pi^{\rm{sup}}}$ of the optimal policy for both the current state $S^{\rm{sup}}_{k}$ and the next state $S^{\rm{sup}}_{k+1}$. Since
$\xi_{k}=A_{\pi^{\rm{sup}}}(S_{k}^{\rm{sup}},a_{k}^{\rm{inf}})|_{a_{k}^{\rm{inf}}\sim\pi^{\rm{inf}}(\cdotp|S_{k}^{\rm{inf}})}=\mathrm{E}_{S^{\rm{sup}}_{k+1}}[\delta_{\pi^{\rm{inf}},k}],$ (32)
the TD error provides an unbiased estimate of the IVoI, assuming the value function estimator accurately approximate the true value function $v_{\pi^{\rm{sup}}}(S^{\rm{sup}}_{k})$. While this method introduces some noise, its variance is relatively low since the expectation in (32) is taken only over the next state $S^{\rm{sup}}_{k+1}$.
This method is particularly attractive when DRL algorithms such as asynchronous advantage actor-critic (A3C) [38] and proximal policy optimization (PPO) [39] are used to learn $\pi^{\rm{sup}}$, as these algorithms inherently estimate the value function while training stochastic policies. However, similar to Method B, the value function estimator may be less accurate for states that are frequently visited by the acting policy $\pi^{\rm{inf}}$ but rarely encountered under the optimal policy $\pi^{\rm{sup}}$. Therefore, fine-tuning the value estimator using a Monte Carlo method — similar to Method A — is recommended.
Figure 4: The IVoMI for the vehicle-following control problem in Case 6 during a specific test episode in the Stop-and-Go scenario, shown alongside the acceleration trajectory of the predecessor ($acc_{i-1,k}$), the position error ($e_{pi,k}$), velocity error ($e_{vi,k}$), and the follower’s acceleration ($acc_{i,k}$) under two policies: $\tilde{\pi}_{i}^{*}$ and $\pi_{i}$.
Case 8.
We estimate EVoMI and IVoMI for the vehicle-following control problem as given in Case 6. The velocity trajectory of the preceding vehicle $i-1$ is extracted from the Next Generation Simulation (NGSIM) dataset [40] when training the following vehicle. The optimal policy $\tilde{\pi}_{i}^{*}$ for the augmented-state SSDP in Case 4 is learned by the Twin Delayed Deep Deterministic policy gradient (TD3) algorithm [41]. This policy is also used as the acting policy $\pi_{i}$ when the predecessor’s acceleration is not available. In this case, $acc_{i-1,k}$ is set to 00 in the state.
The EVoMI is $-3.12$, which is calculated by evaluating the policy over $100$ test episodes in which $acc_{i-1,k}$ is either consistently unavailable or consistently available throughout the entire episode, and then computing the difference in the average total reward between the two settings.
Fig. 4 shows the IVoMI during a specific test episode in the Stop-and-Go scenario, along with the acceleration trajectory of the predecessor ($acc_{i-1,k}$), the position error ($e_{pi,k}$), velocity error ($e_{vi,k}$), and follower’s acceleration ($acc_{i,k}$) under the two policies: $\tilde{\pi}_{i}^{*}$ and $\pi_{i}$.
At each control interval $k$, two actions are selected: one by the optimal policy $\tilde{\pi}_{i}^{*}$ when $acc_{i-1,k}$ is available, and the other by the acting policy $\pi_{i}$ when $acc_{i-1,k}$ is not available, respectively. The action prescribed by the acting policy is executed to determine the next state. Meanwhile, the advantage function — evaluated by the trained estimator using the full state, including $acc_{i-1,k}$, and the action selected by the acting policy $\pi_{i}$ as inputs — is used to quantify the IVoMI.
Note that the EVoMI is a constant negative value as it does not depend on specific state action pairs. This negative value indicates that transmitting the predecessor’s acceleration to the follower is beneficial for vehicle control. In contrast, the IVoI varies across control intervals since it depends on the state action pairs. When $acc_{i-1,k}=0$, IVoMI is zero, indicating the information is not useful in this case. Otherwise, the IVoMI takes on a negative value.
@@ VII-B3 Across Multiple Control Intervals - When and What to Communicate
  • Every Communication Interval - How to Communicate: The communication agent makes RRM decisions at each communication interval $(k,t)$. To enhance spectrum utilization, we consider a setting where $L$ V2V links share the uplink resources orthogonally allocated to $M$ V2I links. With a slight abuse of notation, each V2V link, connecting a predecessor $i-1$ and its corresponding follower $i$, is indexed by $i-1\in\mathcal{L}=\{1,3,5,\ldots,2L-1\}$. Without loss of generality, we assume that every V2I link $m\in\mathcal{M}=\{0,\cdots,M-1\}$ is pre-assigned sub-channel $m$ and transmits with a constant power $P^{\rm I}_{m}$. Moreover, one or more V2V links can reuse the sub-channels of the V2I links for CAM transmission. We use the binary allocation indicator $\theta_{i-1,m,(k,t)}\in\{0,1\}$ to indicate whether V2V link $i-1$ occupies sub-channel $m$ at communication interval $(k,t)$ or not. Moreover, we consider that each V2V link $i-1$ occupies at most one sub-channel, i.e., $\sum_{m=0}^{M-1}\theta_{i-1,m,(k,t)}\leq 1$. Let $P^{\rm V}_{i-1,m,(k,t)}$ denote the transmit power of V2V link $i-1$ over the sub-channel $m$ at communication interval $(k,t)$. Therefore, each communication agent $i-1$ has to make the subchannel allocation $\{\theta_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ and power allocation $\{P^{\rm V}_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ decisions at each communication interval $(k,t)$.
  • Every Control Interval - When to Communicate: Let $\phi_{i-1,k}\in\{0,1\}$ be the binary transmission indicator to indicate whether the sampled CAM $acc_{i-1,k}$ will be transmitted to the follower $i$ or not. At each control interval $k$, each communication agent $i-1$ can make decisions on $\phi_{i-1,k}$.
  • Across Multiple Control Intervals - When to Communicate: Each communication agent $i-1$ makes a one-time decision on whether it will transmit its acceleration or not to its follower $i$. Once the decision is made, the predecessor will always transmit $acc_{i-1,k}$ or not at each control interval $k$.